# Jon Danielsson London School of Economics Andreas Uthemann

### v4

@@ 2.3 System responses
A key challenge for AI working for the macroprudential, but not generally for the microprudential authorities, relates to the dynamic interaction of the financial system participants. A helpful framework for understanding the problem is the Lucas (1976) critique, which states that the decision rules used by economic agents depend on the underlying economic environment and can change as regulations change, undermining their effectiveness. An example of Goodhart’s (1974) law, “Any observed statistical regularity will tend to collapse once pressure is placed upon it for control purposes.” Changes to financial regulations or the level of supervision, including changes to the crisis resolution playbook, as well as decisions taken during resolution, will change the responses of the private sector to a similar crisis in the future in unexpected ways.
Of particular interest when AI is to be used for economic analysis relates to how Lucas and Goodhart came to their conclusions. Prior to the 1970s, economists believed in large general models informed by statistical patterns, with the Phillips curve, representing a trade-off between unemployment and inflation based on a negative empirical correlation of these two variables, as a centrepiece. [^2] For an important example, see Rasche and Shapiro (1968) explaining the MIT-Fed model used by the Federal Reserve and developed in collaboration with MIT. The contribution of Lucas was to recognise that when the authorities try to exploit this apparent trade-off, they change the way the private sector negotiates wages and sets prices, leading to a breakdown of previously identified correlations. Further attempts to stimulate the economy through this channel create inflation without reducing unemployment, giving rise to the word stagflation. In response, the economic profession moved away from large-scale simultaneous equation models based on statistical correlations to smaller, structural models that explicitly modelled beliefs and general equilibrium effects so that reactions to policy interventions followed a well understood and internally consistent logic encoded in the model, perhaps an early example of interpretability. An AI engine would need to do the same if it were to become an effective macropru regulator.
@@
One direct impact is on measuring financial risk, an essential task for any regulator. Besides the sampling issues discussed in Section 2.1 above, there are particular technical issues for why risk measurements can be misleading. It is helpful to use the classification scheme proposed by Daníelsson and Shin (2002), which separates financial risk into two categories: exogenous and endogenous. Exogenous risk emphasises risk measured by statistical techniques based on historical outcomes in financial markets, typically prices. Endogenous risk, by contrast, captures risk that arises from the strategic interaction of the economic agents that make up the financial system.
Exogenous risk is easy to measure, and AI excels at it. Identifying endogenous risk is difficult because it captures outcomes only visible in extreme stress when self preservation and mistrust of counterparties are crisis amplifiers.
@@
The relative importance of exogenous versus endogenous risk depends on the problem. For most microprudential regulations, the frequency of events and lack of large strategic interactions means that assuming risk is exogenous is usually quite innocuous. However, any authority using data driven analytics that uses exogenous risk measurement for assessing the risk of financial instability, will likely be seriously misled as to how the financial system evolves in times of stress.
@@ 2.4 Pre-specified objectives and distributed decision making
The fourth conceptual challenge arises from the clarity of the objectives AI optimises for. In the best case scenario, it knows the objectives, and can in many cases use reinforcement learning to identify solutions in real time. While AI can operate in an environment with mutable objectives, it is less effective and more prone to mistakes as the rarity of events and the cost of mistakes increase. The worst case for AI is when the objective is unknown ex-ante and cannot be learned, and that is where its potential for making catastrophically wrong advice and decisions is the strongest.
@@
Mutable objectives do not pose much of a problem to micropru. The rulebook is known and usually static on the timescale decisions are made. While it evolves in response to events and the regulated response to regulations, AI can quickly update its understanding of the objectives, by adopting changes to the rulebook, learning from observing how the human supervisors act and reinforcement learning.
@@
This is not the case in macropru, as it operates on very long timescales. The time between relevant events is usually in decades. It can be very difficult to define the macroprudential objective except at the highest levels of abstraction, such as the prevention of severe dysfunction in key financial markets and especially the failure of systemically important institutions. This applies to the design of macroprudential regulations, macroprudential supervision and crisis resolution. It can be difficult to make a case for the need to allocate significant resources to prevent something that only happens in the distant future, and the macroprudential regulations are susceptible to lobbying and political interference, which means that the objectives of regulations can change over time. Learning from previous crisis interventions might not be of much help as the concrete goals the actions under consideration are to achieve are likely specific to particular circumstances and political environments.
@@
The most severe financial crises can have catastrophic consequences if not addressed adequately, with direct economic costs in the several trillions of dollars, as noted by Barnichon et al. (2022), and a large number of people materially affected. When that happens, society demands we do what it takes to resolve the crisis. Often, the rules and the laws in place might stand in the way of the most effective crisis resolution. Emergency sessions of Parliament to rectify that are not uncommon, such as Switzerland’s resolution of Credit Suisse. [^3] https://www.bloomberg.com/news/articles/2023-03-20/credit-suisse-collapse-reveals-some-ugly-truths-about-switzerland-for-investors. Emergency constitution clauses might be invoked. Pistor (2013), in her legal study of the resolution of financial crises, finds that if the existing law prevents the most effective course of action, there is acceptance from the political and judicial system to suspend the law in the name of the higher objective of crisis resolution. Furthermore, when a severe crisis happens, the political leadership takes charge, inevitable because if it becomes necessary to change or bypass the law or significantly redistribute resources, the political leadership is the only entity with the necessary legitimacy. Given the fluidity of this process, it is difficult to see how one would specify the objectives for an AI so that it can provide real-time advice for crisis resolution or make decisions.
@@
We have a long experience of resolving crises and have a relatively good understanding of the process. The regulatory system is usually modular, with separate authorities and fiercely guarded mandates. In the most severe crises, these silos break down. All relevant authorities, the affected private sector, the judiciary, and especially the political leadership have come together to decide how to resolve the crisis. Government ministers usually lead this process. This may involve the same entities in other financial centres in a global crisis. Each stakeholder brings their own education, ethics, professional experience and objectives to the table. Such a process can be highly robust. All pertinent issues are discussed, including information that was confidential or implicit until then. Such analysis depends on implicit knowledge and intuitive understanding its participants have of the current situation and each others’ views and objectives. Intuition can mean finding analogous problems from seemingly unrelated domains that can provide creative solutions for the current crisis that has not been encountered before. This process of successful extrapolation is often vital to crisis resolution and can imply ignoring or revaluing previously pursued goals. There are many such examples in history. In his re-evaluation of the 1866 crisis, Bagehot (1873) invented modern financial stability policies, like lending of last resort, while the German central banker, Hjalmar Schacht, used a short squeezing to stop the hyperinflation in 1923. The Swedish central bank in 1992 created the good bank-bad bank model for crisis resolution. Reproducing this process will be particularly difficult for AI as it implies both shifting attention to seemingly unrelated problems and potentially modifying its objective function.
@@
Crisis resolution is arguably the most important aspect of financial policy, especially for central banks — one of their raisons d’être. Nevertheless, given the above challenges — scarce data, unknown-unknowns, endogenous structural changes in response to attempted control — humans struggle with this task, and this is where AI could benefit them the most. Unfortunately, it has to overcome the same conceptual problems, and this challenge is as difficult for AI as for human regulators, if not more so. Paradoxically, progress in AI’s suitability for the task might come from a better human understanding of these problems that can then be translated into better algorithms.
@@ 2.5 Incentives
The objectives of the various market participants are specific to them and often in conflict with those of other participants and society. That is why we have financial regulations: to align the private sector’s incentives with society at large. The interactions between financial supervisors and the private sector decision-makers can be seen as a principal-agent (PA) relationship, where the highly resourced and sophisticated private sector — the agent — wants to get on with successfully running its business. In contrast, the supervisors — the principal — aim to ensure that the agents operate a financial system that efficiently channels savings to investments while not abusing their clients or taking excessive risks that can cause costly crises.
@@
The PA problem applied to financial markets is particularly difficult for three reasons. First, contracting is incomplete, meaning we cannot specify what will happen in every contingency. We might not even know what the contingencies are. Second, the very behaviour to be controlled is risk taking, and since risk is a latent variable, it cannot be measured directly but only imprecisely inferred from market movements. Consequently, it can be difficult to determine whether an undesirable outcome is due to misbehaviour or innocent bad luck. Finally, because implicit or explicit government guarantees are fundamental to all financial regulations, market participants, knowing that some of their risk taking is publicly insured, can behave in a way that makes government bailouts more likely, creating moral hazard. [^4] Dewatripont et al. (1994) provides a detailed analysis of the problem of financial regulation with incomplete contracts, government guarantees and private sector moral hazard.
@@
As we increasingly use AI in regulations and the private sector, a new dimension in the PA problem emerges. The one-sided financial institution-regulator problem becomes two-sided, with the addition of the relationship between the regulators and financial institutions as principals and their respective AI as agents. The reason is that we cannot expect the AI engines’ objectives to be fully aligned with its human operator as AI can take hidden actions to achieve its goals. This means the human operator needs to incentivise the AI to act in its owner’s interest.
The presence of the two-sided PA problem amplifies the already significant existing PA problem in financial regulation. We have in the past seen many cases where market participants actively and deliberately create heightened stress since anybody forewarned of stress can profit. Creating and amplifying crises is profitable. Lowenstein (2000) shows just one example from the LTCM crisis in 1998. AI might find it even easier to act in that way. It will, therefore, be essential for the authorities to explicitly consider how to align the AI engines, whether directly operated by the regulated or by their AI vendors, with the objectives of the authority.
@@
Existing mechanisms for managing the regulator-bank PA relationship will not be adequate when AI is involved. In the technical terminology of moral hazard, AI has more hidden actions at its disposal and will command higher information rents. In addition, AI’s objectives can be less straightforward than those of traditional financial institutions that trade-off risks and returns. How do you incentivise an algorithm whose objective is to predict the next token correctly? What higher level objectives does this loss function induce, and how do we formulate an incentive scheme that imposes the correct constraints on AI behaviour? How can we make it internalise negative externalities, such as the possibility of deliberately creating fire sales or runs to trigger privately beneficial but publicly costly bailouts?
3 Artificial intelligence and machine learning
@@
There is no standard definition of artificial intelligence (AI). One of the pioneers of AI, John McCarthy, saw it as “the science and engineering of making intelligent machines”, [^5] http://jmc.stanford.edu/artificial-intelligence/what-is-ai/index.html where perhaps “intelligence measures an agent’s ability to achieve goals in a wide range of environments” in the definition of Legg and Hutter (2007). Generally, it is a computer algorithm that performs tasks that a human would otherwise do. For a general view of AI, Russel (2019) is particularly useful.
AI distinguishes itself from machine learning (ML) and traditional statistical modelling in that it is also used to do analysis, provide recommendations and make decisions, not merely give quantitative inputs for human decision making as a statistical algorithm would. AI charged with tasks aims to find the best course of action given its objectives and its understanding of the problem domain.
ML provides the fundamental building blocks of most AI, finding statistical regularities in large data sets and helping provide AI engines with a deeper understanding of the underlying latent structure that generates the data of a given problem domain. One of the key differences to conventional applied statistics is that the input data tend to be unstructured, are often text-based and large in size, where the models can encode highly complex statistical relationships. The modeller usually does not supply a prior understanding of the parameterised model structure. See the Appendix for details on ML that are useful in financial applications.
@@ 3.1 From reactive machines to AGI
Even though the ability of AI engines has increased rapidly, today’s AI merely match patterns and do not understand in the same way a human does. It is unclear what achieving such an understanding will involve. When considering the abilities of various AIs, including those only hypothesised, it is beneficial to classify AI based on its abilities. As it turns out, that is a surprisingly difficult task since the experts do not agree on classification, terminology, ability or future potential. At the risk of oversimplification, there are four types of AI.
@@
The second category is limited memory AI, which both has knowledge of what happened in the past and keeps track of past actions and observations and can condition its response to a given situation on these past experiences. Generative AI, such as ChatGPT, fall into this category, systems that predict the next word or phrase, or some other object, based on the content it is working with. Other examples include virtual assistants and self driving cars. Time series models are also of this type.
@@
And finally, artificial general intelligence or AGI. It is still purely theoretical while often hypothesised. It would use existing knowledge to learn new things without needing human beings for training and being able to perform any intellectual task humans can. And if it reaches that level, increased computer capacity would allow it to surpass humans. [^6] Kasirzadeh discusses existential risks arising from AGI
@@
A point of controversy is whether the next token prediction is sufficient for achieving AGI. In that case, all that is required for AGI is a sufficient amount of training data and compute. We would not need conceptual improvements in network technology.
When transiting from using AI only to learn about the world to use it to advise on real-world decisions and even make such decisions, a key challenge is AI alignment and safety, that is, how to make AI systems behave in line with human intentions and values. [^7] For a survey of AI alignment work, see, e.g., Ji et al. (2023), and Shevlane et al. (2023), and Bengio et al. (2023) model evaluation for extreme risks.

### v5

@@ 2.3 System responses
A key challenge for AI working on macropru, but not generally for the microprudential authorities, arises from the dynamic interaction of the financial system participants. A helpful framework for understanding the problem is the Lucas (1976) critique, which states that the decision rules used by economic agents depend on the underlying economic environment and can change as regulations change, undermining their effectiveness. Goodhart, in his critique, applies this to regulations, stating, “Any observed statistical regularity will tend to collapse once pressure is placed upon it for control purposes.” Changes to financial regulations or the level of supervision, including changes to the crisis resolution playbook, as well as decisions taken during resolution, will change the responses of the private sector to similar crises in the future in a way that is unexpected.
@@
A particular consequence is the measurement of financial risk, an essential task for any regulator. Besides the data issues discussed above, there are particular technical issues for why risk measurements can be misleading. It is helpful to use the classification scheme proposed by Daníelsson and Shin (2002), which separates financial risk into two categories: exogenous and endogenous. Exogenous risk emphasises risk measured by statistical techniques based on historical outcomes in financial markets, typically prices. Endogenous risk, by contrast, captures risk that arises from the strategic interaction of the economic agents that make up the financial system.
Exogenous risk is easy to measure, and AI excels at it. Identifying endogenous risk is difficult because it captures outcomes only visible in extreme stress when self-preservation and mistrust of counterparties are crisis amplifiers. AI will help identify and cope with some forms of endogenous risk. In times of stress, humans might specify the objectives, while AI might be better at reacting to a crisis given those objectives. Even then, it will have to contend with the socially damaging individual optimal mistrust and self-preservation. Ultimately, an AI engine interacting with another AI engine would likely face similar issues as with humans facing humans today, pre-emptive runs, multiple equilibria and the like. These outcomes happen even in fully rational settings, and AI-advised and managed systems might be more susceptible to such outcomes than more human-centred arrangements.
@@
The relative importance of exogenous versus endogenous risk depends on the problem. For most microprudential regulations, the frequency of events and lack of large strategic interactions mean that assuming risk is exogenous is usually quite innocuous. However, any authority using data-driven analytics that uses exogenous risk measurement to assess the risk of financial instability will likely be seriously misled as to how the financial system evolves in times of stress, as noted by Daníelsson et al. (2017).
@@ 2.4 Pre-specified objectives and distributed decision making
While both humans and AI can operate in an environment with mutable objectives, both become less effective and more prone to mistakes as events become infrequent. The worst case is when the objective is unknown ex-ante and cannot be learned, when its potential for providing seriously flawed advice and catastrophically bad decisions is the strongest.
@@
Mutable objectives do not pose much of a problem to micropru. The rulebook is known, and usually static on the timescale decisions are made. While it evolves in response to events and the response of the regulated to regulations, AI can quickly update its understanding of the objectives by adopting changes to the rulebook, by observing how the human supervisors act.
@@
This is different in macropru, as it operates on very long timescales. The time between relevant events is usually in decades. It can be very difficult to define the macroprudential objective except at the highest levels of abstraction, such as the prevention of severe dysfunction in key financial markets and especially the failure of systemically important institutions. This applies to the design of regulations, supervision and crisis resolution. It is not easy to make a case for the need to allocate significant resources to prevent something that only happens in the distant future. Meanwhile, macroprudential regulations are susceptible to lobbying and political interference, which means that the objectives of regulations change over time. While it is straightforward to set the policy objective to “prevent the failure of SIFIs”, the actual implementation of such a policy has widespread implications on the entire financial system, and the authorities making such decisions in real-time will draw on facts and interests only visible at the time. Learning from previous crisis interventions might be of little help as the concrete goals the actions under consideration are to achieve are likely specific to particular circumstances and political environments.
The most severe financial crises can have catastrophic consequences if not addressed adequately, with direct economic costs in the several trillions of dollars, as noted by Barnichon et al. (2022), with a large number of people materially affected, leading to considerable political consequences as well. When such crises happen, society demands we do what it takes to resolve the crisis. Chwieroth and Walter (2019) note that bailouts have become what they call a middle-class good, making them inevitable.
@@
Often, the rules and the laws in place stand in the way of the most effective crisis resolution. Emergency sessions of Parliament to rectify that are not uncommon, such as Switzerland’s resolution of Credit Suisse. [^4] https://www.bloomberg.com/news/articles/2023-03-20/credit-suisse-collapse-reveals-some-ugly-truths-about-switzerland-for-investors Emergency constitution clauses might be invoked. Pistor (2013), in her legal study of the resolution of financial crises, finds that if the existing law prevents the most effective course of action, there is acceptance from the political and judicial system to suspend the law in the name of the higher objective of crisis resolution. Furthermore, when a severe crisis happens, the political leadership takes charge, inevitably because if it becomes necessary to change or bypass the law or significantly redistribute resources, the political leadership is the only entity with the necessary legitimacy.
@@
Given the fluidity of this process, will AI find it particularly difficult to provide real-time advice? We have a long experience in resolving crises and a relatively good understanding of the process. The regulatory system is usually modular, with separate authorities and fiercely guarded mandates. In the most severe crises, these silos break down. All relevant authorities, the affected private sector, the judiciary, and especially the political leadership have come together to decide how to resolve the crisis. Government ministers usually lead this process. This may involve the same entities in other financial centres in a global crisis. Each stakeholder brings their own education, ethics, professional experience and objectives to the table. Such a process can be highly robust. All pertinent issues are discussed, including information that was confidential or implicit until then. Such analysis depends on implicit knowledge and intuitive understanding its participants have of the current situation and each others’ views and objectives. Intuition can mean finding analogous problems from seemingly unrelated domains that can provide creative solutions for the current crisis that has not been encountered before. This process of successful extrapolation is often vital to crisis resolution and can imply ignoring or revaluing previously pursued goals. There are many such examples in history. The German central banker Hjalmar Schacht used short squeezing to stop the hyperinflation in 1923. The Swedish central bank in 1992 created the good bank-bad bank model for crisis resolution. Suppose AI is to become an effective macropru regulator. In that case, it will have to reproduce such a process and pay attention to nuance, hidden objectives and mutable objectives.
@@
Crisis resolution is arguably the most important aspect of financial policy, especially for central banks — one of their raisons d’être. Nevertheless, given the above challenges — scarce data, unknown-unknowns, endogenous structural changes in response to attempted control — humans struggle with this task, and this is where AI could benefit them the most. Unfortunately, it has to overcome the same conceptual problems, and this challenge is as difficult for AI as for human regulators, if not more so.
Paradoxically, progress in AI’s suitability for the task might come from a better human understanding of these problems that can then be translated into better algorithms.
@@ 2.5 Incentives
The objectives of the various market participants are specific to them and often in conflict with those of other participants and society. In the past, we have seen cases where market participants actively and deliberately create heightened stress since anybody forewarned of stress can profit. Creating and amplifying crises is profitable. Lowenstein (2000) shows just one example from the LTCM crisis in 1998, and what is more frequent, their individually optimal actions — run for safety, survival — endogenously amplifies socially undesirable stress. That is why we have financial regulations: to align the private sector’s incentives with society. The interactions between financial supervisors and the private sector decision-makers can be seen as a principal-agent (PA) relationship, where the highly resourced and sophisticated private sector — the agents — wants to get on with successfully running their businesses. In contrast, the supervisors — the principals — aim to ensure that the agents operate a financial system that efficiently channels savings to investments while not abusing their clients too much or taking excessive risks that can cause costly crises.
@@
The PA problem applied to financial markets is particularly difficult for three reasons. First, contracting is incomplete, meaning we cannot specify what will happen in every possible contingency. We might not even know what the contingencies are. Especially as the very behaviour to be controlled, risk-taking, means there are circumstances that the principal does not anticipate. Risk is a latent variable that cannot be measured directly but only imprecisely inferred from market movements. Consequently, it can be difficult to determine whether an undesirable outcome is due to misbehaviour or innocent bad luck. Second, because implicit or explicit government guarantees are fundamental to all financial regulations, market participants, knowing that some of their risk-taking is publicly insured, can behave in a way that makes government bailouts more likely, creating moral hazard. [^5] Dewatripont et al. (1994) provides a detailed analysis of the problem of financial regulation with incomplete contracts, government guarantees and private sector moral hazard.
@@
As we increasingly use AI in regulations and the private sector, a new dimension in the PA problem emerges. The one-sided financial institution-regulator problem becomes two-sided, with the addition of the relationship between the regulators and financial institutions as principals and their respective AI as agents. The reason we worry about the AI’s hidden actions is that objectives are misaligned. This means the human operator of AI needs to incentivise its AI to act in humans’ interest, so there is also a principal-agent problem between the agent and its AI.
@@
The presence of the two-sided PA problem amplifies the already significant existing PA problems in financial regulation. The authorities will have to explicitly consider how to align the AI engines, whether directly operated by the regulated or by their AI vendors, with the objectives of the authority. Existing mechanisms for managing the regulator-bank PA relationship will not be adequate when AI is involved. In the technical terminology of moral hazard, AI has more hidden actions at its disposal and will command higher information rents. While AI’s objectives might be simpler, its actions/strategies to achieve the objectives may be more complex. It leads to difficult questions. What higher-level objectives does this loss function induce, and how do we formulate an incentive scheme that imposes the correct constraints on AI behaviour? How can we make it internalise negative externalities, such as the possibility of deliberately creating fire sales or runs to trigger privately beneficial but publicly costly bailouts?
3 Artificial intelligence
@@
There is no standard definition of artificial intelligence (AI). One of the pioneers of AI, John McCarthy, saw it as “the science and engineering of making intelligent machines”666jmc.stanford.edu/artificial-intelligence/what-is-ai/index.html where perhaps “intelligence measures an agent’s ability to achieve goals in a wide range of environments” in the definition of Legg and Hutter (2007). ChatGPT defines it as “The capability of a machine to imitate intelligent human behaviour, including learning, reasoning, problem-solving, perception, and language understanding”. We see it as a computer algorithm performing tasks a human would otherwise do, (Norvig and Russell, 2021). In Russell’s (2019) taxonomy, AI as a rational maximising agent resonates with the economic notion of utility maximising agents and hence is particularly helpful in the analysis of AI use in the financial system.
While statistical and machine learning [^7] ML provides the fundamental building blocks of most AI, finding statistical regularities in large data sets and helping provide AI engines with a deeper understanding of the underlying latent structure that generates the data of a given problem domain. See the Appendix for details on ML that are useful in financial applications. (ML) algorithms are limited to giving quantitative inputs for humans making decisions; AI also provides recommendations and even makes decisions. When we charge AI with a task, it searches for the best course of action given its objectives and understanding of the problem domain. A key differentiator between the conventional and AI way of operating is how AI learns in real-time while interacting with the environment.
@@ 3.1 From reactive machines to AGI
Even though the ability of AI engines has increased rapidly, today’s AI merely matches patterns and does not understand them in the same way as humans. It is unclear what achieving such an understanding will involve. When considering the abilities of various AIs, including those only hypothesised, it is beneficial to classify AI based on its abilities. That is a surprisingly difficult task since the experts do not agree on classification, terminology, ability or future potential. At the risk of oversimplification, there are four types of AI.
@@
The second category is limited memory AI, which knows what happened in the past, keeps track of past actions and observations and can condition its response to a given situation on these past experiences. Generative AI, such as ChatGPT, falls into this category, as it is a system that predicts the next word or phrase or some other object based on the content it is working with. Other examples include virtual assistants and self-driving cars. Time series models are also of this type.
@@
And finally, artificial general intelligence or AGI, which is still purely theoretical while often hypothesised. It would use existing knowledge to learn new things without needing human beings for training and being able to perform any intellectual task humans can. If it reaches that level, increased computer capacity will allow it to surpass humans.
@@
A further point of controversy is whether the next token prediction is sufficient for achieving AGI, since in that case, all that is required for AGI is a sufficient amount of training data and compute, conceptual improvements in network technology are not needed.